ahmer walks 100m east 70m north 140m east find the displacement and distance?
1
Expert's answer
2020-01-14T09:28:45-0500

The distance


"d=d_1+d_2+d_3=100+70+140=310\:\rm m"

The displacement


"{\bf s}={\bf d}_1+{\bf d}_2+{\bf d}_3"

where


"{\bf d}_1=100{\bf i}+0{\bf j},\; {\bf d}_2=0{\bf i}+70{\bf j}, \; {\bf d}_3=140{\bf i}+0{\bf j}"

Hence


"{\bf s}=100{\bf i}+0{\bf j}+0{\bf i}+70{\bf j}+140{\bf i}+0{\bf j}=240{\bf i}+70{\bf j}"

The magnitude of a displacement


"s=\sqrt{(240)^2+(70)^2}=250\:\rm m"